# Cascaded Multi-Scale Attention for Enhanced Multi-Scale Feature Extraction and Interaction with Low-Resolution Images

### v1

@@ Abstract
keywords:
[…]
journal: Computer Vision and Image Understanding
1 Introduction
In recent years, the application of deep learning for image recognition has significantly broadened its scope in the real world. At the forefront, there is the use under tougher conditions for image recognition. One of these is the application to low-resolution images. In this paper, we discuss how to maximize the estimation accuracy for several typical image recognition problems using images of a lower resolution than those primarily considered in previous studies, as illustrated in the left panel of Figure  1.
Fig. 1: Left: Estimating pose from low-resolution images. Left-Upper: human pose estimation from COCO2017. Left-Lower: head pose estimation from AFLW2000. Right: resolution-accuracy trade-off for human pose estimation evaluated on COCO 2017 val dataset.
@@ 1 Introduction
Why do we aim to do this? First, multi-scale feature extraction combined with their interactions is an essential component for several tasks. The estimation of human poses mentioned earlier is a prime example of this necessity, as evidenced by previous studies (Cheng et al., 2020; Yuan et al., 2021b; Wang et al., 2022c). While the level of necessity may differ, this strategy is fundamentally beneficial across a broad spectrum of image recognition tasks, including, for example, fine-grained object recognition (Ibh et al., 2023) and biomedical image segmentation (Heidari et al., 2023).
@@
Second, low resolution complicates the management of multi-scale features. The conventional approach generates multi-scale features through the downsampling of feature maps (Sun et al., 2019; Cheng et al., 2020; Yuan et al., 2021b; Yu et al., 2021). This method is effective when starting with high-resolution images. However, with low-resolution inputs, repeated downsampling quickly diminishes the resolution of feature maps to the point where additional downsampling results in a significant loss of information.
@@
To address this challenge, we investigate methods to manage multi-scale feature maps without downsampling, enabling interactions between features of varying scales. Specifically, we first divide the multi-heads in Multi-Head Self-Attention (MHSA) into several groups, each of which processes features of a different scale. Then, instead of downsampling, we adopt the window-based self-attention mechanism from the Swin Transformer (Liu et al., 2021) to extract a different scale feature in each group. To facilitate effective interactions among features of different scales, we transfer a feature map from one group to another in an order from lower to higher scales. The transferred feature is utilized in the attention calculations of the subsequent group.
@@
The approach to handling multi-scale features through grouping heads in the first part, and the cascading calculation structure in the latter part, are inspired by recent CNN-ViT hybrid models, SG-Former (Ren et al., 2023) and EfficientViT (Liu et al., 2023), respectively. It is important to note that these existing models, while aiming to address the high computational demands of the ViT architecture and improve the trade-off between performance and computation, were not designed with low-resolution images or the extraction of multi-scale features plus their interactions in mind.
@@
We present experimental results across three distinct tasks: human pose estimation, head pose estimation, and image classification, all utilizing low-resolution image inputs. The results demonstrate that our proposed method achieves superior performance than existing state-of-the-art methods tailored for these individual tasks; see the right panel of Figure  1. It is also noted that our method needs fewer parameters than the existing methods.
2 Related work
2.1 Extraction and Use of Multi-Scale Features
@@
Integrating features from multiple scales improves neural networks’ ability to deeply understand complex visual scenes. HRNet (Sun et al., 2019; Cheng et al., 2020) and HRFormer (Yuan et al., 2021b) fuse the multi-scale representations from multiple branches to effectively generate reliable high-resolution representations with strong position sensitivity. CrossViT (Chen et al., 2021) learns multi-scale feature representations for image recognition by a dual-branch transformer and merges them multiple times to complement each other. TopFormer (Zhang et al., 2022) produces scale-aware global features from tokens at various scales. P2T (Wu et al., 2022) and Shunted Transformer (Ren et al., 2022) improve upon PVT (Wang et al., 2021) by applying various downsampling rates for the keys and values across heads, rather than using identical spatial reduction for both. This allows each head to capture both coarse- and fine-grained details. SG-Former (Ren et al., 2023) aggregates tokens at multiple scales and performs attention mechanisms to extract global and fine-grained information for visual tasks. SMT (Lin et al., 2023b) captures diverse spatial features through multiple convolutions with different kernel sizes and conducts scale-aware aggregation to enhance information interaction. SBCFormer  (Lu et al., 2024) achieves efficient global and local feature fusion by introducing a bi-branch CNN-ViT hybrid module. In existing methods, multi-scale representations are typically generated by downsampling the feature map, which can be problematic for low-resolution inputs, as it exacerbates the loss of detail. Our study mitigates this issue by enabling multi-scale feature extraction and interaction without the need for downsampling.
2.2 Dealing with Low-Resolution Images
@@
Low-resolution images often lack clarity in fine textures, edges, and small details, presenting significant challenges in image analysis. Some studies tackle this issue by incorporating augmented inputs or utilizing additional knowledge from fully trained teacher models. DeriveNet (Singh et al., 2021) introduces a multi-resolution pyramid-based data augmentation technique to learn effective features from low-resolution images. NLSA (Mei et al., 2021) proposes a non-local sparse attention to explore long-range feature correlations in low-resolution images. CAL (Wang et al., 2022a) employs Gaussian offset weighting to account for displacements between coarse predictions and ground-truth positions in low-resolution images. FMD (Huang et al., 2022) proposes a feature distillation framework to transfer high-definition knowledge from a teach model for improved low-resolution object recognition. On the other hand, in order to reconstruct fine details from low-resolution images, recent Transformer-based approaches are gaining interest. ESRT (Lu et al., 2022) combines a CNN backbone with a Transformer backbone to recover a super-resolution image from its degraded low-resolution counterpart. ESSAformer (Zhang et al., 2023c) proposes an attention-embedded Transformer to enlarge the receptive field, allowing for better extraction of information from low-resolution hyperspectral images. Steformer (Lin et al., 2023a) designs a bi-branch Transformer-based structure to extract and integrate cross-view information from low-resolution stereo images. However, existing methods do not explore feature fusion in multi-scale space to improve low-resolution image recognition. In this paper, we propose a cascaded multi-scale attention mechanism that constructs enriched features in multi-scale spaces for low-resolution images.
2.3 CNN-ViT Hybrids
@@
Hybrid models that combine convolutions and Vision Transformer (ViT) (Dosovitskiy et al., 2021) are gaining interest for their ability to capture fine-grained local details while preserving global dependency. To enhance fine-grained information, some methods apply the attention mechanisms to sub-regions or tokens within feature maps. Swin (Liu et al., 2021) employs shifted windows in the self-attention mechanism to capture both local and global information. CSwin (Dong et al., 2022) introduces a cross-shaped window to enhance the self-attention mechanism and further improve model capacity. NAT (Hassani et al., 2023) proposes a scalable sliding window attention mechanism by localizing self-attention to the nearest neighboring pixels. QFormer (Zhang et al., 2024) extends the window-based attention mechanism within adaptive quadrangles for learning better feature representation. With the demand for facilitating the aggregation of spatial information, many studies integrate convolutions within the attention mechanism. MobileViT (Mehta and Rastegari, 2022) and EfficientFormer (Li et al., 2022b) combine convolutions with Transformers to achieve efficient local and global feature extraction. CvT (Wu et al., 2021), IRMB (Zhang et al., 2023b) and SBCFormer (Lu et al., 2024) introduce depth-wise separable convolutions into the self-attention mechanism to improve representational power. Recent CNN-ViT hybrid methods demonstrate promising performance in high-resolution image recognition, but their accuracy is limited by the inherent challenges of low-resolution images. In this paper, we propose an effective hybrid network architecture tailored for low-resolution images.
3 Attention Mechanism for Low-Resolution Inputs
3.1 Motivation
@@
3.2 Basic Ideas: Cascaded Multi-Scale Attention (CMSA)
There are two primary ideas. The first involves directing each head of the multi-head self-attention mechanism to handle features from different scales. While conventional multi-head self-attention mechanisms already possess sufficient flexibility, they may not inherently fulfill this role. Contrarily, previous research has shown that attention maps tend to exhibit similarity between heads, leading to redundancy (Heo et al., 2021; Bolya et al., 2022; Chen et al., 2022).
To address this, we propose organizing the multi-heads into groups and implementing SwinT’s local window attention (Liu et al., 2021) within each group; see Figure 2c. We vary the window size for each group. This strategy aims to force different attention heads to effectively manage features of different scales. Our approach draws inspiration from the hybrid-scale attention (HSA) in SG-Former (Ren et al., 2023). However, HSA does not use SwinT’s local attention and instead adjusts key and value through token merging [^2] While we initially experimented with the hybrid-scale attention mechanism, we found that our method based on local attention works much better..
Fig. 2: Design of the proposed model. (a) The overall architecture. (b) Design of the basic blocks. (c) Proposed cascaded multi-scale attention (CMSA). See texts for details.
Our second idea involves implementing a new mechanism to enhance interaction among features of different scales across various groups. This entails organizing the groups in a specific order and transferring features obtained from local attention within one group to the next. These features are subsequently integrated into the attention computation within each group, resulting in an architecture that operates in a cascade-style fashion, as depicted in Figure  2c.
This approach draws inspiration from EfficientViT (Liu et al., 2023). It divides multi-head attention into groups and performs cascaded information integration, and our method follows the same principle. However, in EfficientViT, processing within groups primarily employs standard self-attention, and the scale of feature maps is the same across all groups. Its aim is to improve the trade-off between computational cost and expressive power. In contrast, our approach emphasizes fostering interactions between features of different scales. We also design a new method for integrating information between groups to achieve this goal.
Table 1: CMSA variants with different model sizes.
@@ 3.2 Basic Ideas: Cascaded Multi-Scale Attention (CMSA)
Stem 32$\times$24 Patch Embed. $\times$ 2 ($k=3\times 3$)
@@
1 32$\times$24 CMSA Block $\times$ 2 $\times$ 2 $\times$ 2
@@
$head$=1
2 16$\times$12 Patch Embed. $\times$ 1 ($k=3\times 3$)
@@
$head$=3
3 8$\times$6 Patch Embed. $\times$ 1 ($k=3\times 3$)
@@
$head$=4
3.3 Details of CMSA
We will now describe the cascaded multi-scale attention mechanism; see Figure  2c. Let $\mathbf{X}$ be the input to our cascaded multi-scale attention mechanism; $\mathbf{X}$ is a feature map with the size ${H\times W\times C}$. We first apply point-wise convolutions (PW-Conv) to $\mathbf{X}$ to obtain $\mathbf{Q}\in\mathbb{R}^{H\times W\times D}$, $\mathbf{K}\in\mathbb{R}^{H\times W\times D}$, and $\mathbf{V}\in\mathbb{R}^{H\times W\times D}$.
@@ 3.3 Details of CMSA
To deal with multi-scale features, we employ a group structure, as mentioned earlier. Specifically, we divide the heads into $n$ groups, enabling each group to extract and foster features with different scales. Thus, $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are segmented in their channel dimensions to yield query, key, and value for each group as $\mathbf{K}_{1}$,…,$\mathbf{K}_{n}$, $\mathbf{Q}_{1}$,…,$\mathbf{Q}_{n}$, and $\mathbf{V}_{1}$,…,$\mathbf{V}_{n}$; $\mathbf{Q}_{k}$, $\mathbf{K}_{k}$, $\mathbf{V}_{k}\in\mathbb{R}^{H\times W\times d_{k}}$.
@@
So far, we have considered the standard setting of attention where the query, key, and value all share the same dimensions. Prior studies (Graham et al., 2021; Wang et al., 2022b; Guo et al., 2022) demonstrate that making the key and value spatially smaller (or equivalently, using fewer tokens) than the query can enhance computational efficiency with the minimum sacrifice of inference accuracy. Specifically, we adjust the convolution in the spatial fusion (SF) module to halve the spatial dimensions of the key and value compared to the query, while doubling the channel size of the value relative to both the query and key. In our experiments, we employ this method to achieve a better trade-off of accuracy and computational cost.
3.4 Design of a Whole Network
3.4.1 Overall Design
We adopt a hierarchical pyramidal architecture, in line with recent CNN-ViT hybrids; see Figure  2a. The architecture starts with a ‘Stem’ block, followed by a sequence of stages. Each stage comprises multiple blocks that include the cascaded multi-scale attention mechanism explained above. A patch embedding layer is placed between stages to reduce the spatial resolution of feature maps.
@@ 3.4.1 Overall Design
Table 2: Results on bottom-up human pose estimation with the COCO 2017 val dataset. means the values reported in the respective papers. * indicates models trained with a different learning rate scheduler.
Model Backbone Input Params COCO 2017 val
[…]
SimpleBaseline (Xiao et al., 2018) ResNet-50 $128\times 96$ 34.0 59.3 85.5 67.4 57.8 63.8 66.6
HRNet-W32 (Sun et al., 2019) HRNet-W32 $128\times 96$ 28.5 66.9 88.7 76.3 64.6 72.3 73.7
HRNet-W48 (Sun et al., 2019) HRNet-W48 $128\times 96$ 63.6 68.0 88.9 77.4 65.7 73.7 74.7
DARKZhang et al. (2020a) HRNet-W32 $128\times 96$ 28.5 70.7 88.9 78.4 67.9 76.6 76.7
UDP (Huang et al., 2020) HRNet-W32 $128\times 96$ 28.6 69.5 91.3 77.7 67.9 72.2 73.7
HRFormer-B  (Yuan et al., 2021b) HRFormer-B $128\times 96$ 43.0 66.2 88.2 76.1 64.1 71.8 73.0
ViTPose-B  (Xu et al., 2022) ViT-B $128\times 96$ 86.0 67.8 88.3 76.2 65.9 73.3 74.6
PCT-B (Geng et al., 2023) Swin-B $128\times 96$ 87.0 66.4 90.5 75.7 65.1 68.7 70.1
CMSA-B CMSA-B $128\times 96$ 5.6 70.3 91.3 77.8 67.9 77.4 74.6
CMSA-L CMSA-L $128\times 96$ 7.3 71.5 91.4 79.2 69.0 75.4 75.5
HRNet-W32 (Sun et al., 2019) HRNet-W32 $64\times 48$ 28.5 58.3 86.2 67.1 57.4 62.3 66.2
CAL  (Wang et al., 2022a) HRNet-W48 $64\times 48$ 110.3 61.5 88.1 68.7 60.7 63.5 64.8
UDP (Huang et al., 2020) HRNet-W32 $64\times 48$ 28.6 62.8 86.4 70.9 61.3 67.8 70.4
HRFormer-B  (Yuan et al., 2021b) HRFormer-B $64\times 48$ 43.0 60.9 87.1 70.6 60.0 65.2 68.7
ViTPose-B  (Xu et al., 2022) ViT-B $64\times 48$ 86.0 63.8 86.5 72.6 62.5 68.7 71.4
PCT-B (Geng et al., 2023) Swin-B $64\times 48$ 87.0 62.0 90.5 71.9 61.4 63.8 66.7
CMSA-B CMSA-B $64\times 48$ 5.6 64.7 87.6 72.3 62.9 69.7 71.7
CMSA-L CMSA-L $64\times 48$ 7.5 65.9 88.1 74.5 64.0 70.2 72.7
CMSA-B CMSA-B $64\times 48$ 5.6 65.2 91.4 75.3 63.0 68.7 69.3
CMSA-L CMSA-L $64\times 48$ 7.3 66.0 91.5 76.4 63.6 69.5 70.0
HRNet-W32  (Sun et al., 2019) HRNet-W32 $32\times 24$ 28.5 42.6 79.6 41.5 43.4 44.6 52.7
UDP (Huang et al., 2020) HRNet-W32 $32\times 24$ 28.6 51.9 81.7 56.8 51.5 55.9 61.4
HRFormer-B  (Yuan et al., 2021b) HRFormer-B $32\times 24$ 43.0 45.4 81.8 46.2 45.9 48.0 55.3
ViTPose-B  (Xu et al., 2022) ViT-B $32\times 24$ 86.0 51.2 81.0 56.1 50.7 55.0 60.4
PCT-B (Geng et al., 2023) Swin-B $32\times 24$ 87.0 44.4 84.6 41.8 45.3 43.7 50.7
CMSA-S CMSA-S $32\times 24$ 4.1 52.3 84.0 57.3 51.6 53.9 57.2
CMSA-B CMSA-B $32\times 24$ 5.6 53.5 84.9 58.1 53.2 54.7 58.5
CMSA-L CMSA-L $32\times 24$ 7.3 56.4 87.0 62.6 55.6 58.2 61.1
3.4.2 Design of a Block
As shown in Figure  2b, a single block consists of three components: position embedding, the cascaded multi-scale attention, and a feed-forward network.

### v3

@@ Abstract
Index Terms:
Article submission, IEEE, IEEEtran, journal, , paper, template, typesetting.
I Introduction
In recent years, the application of deep learning for image recognition has significantly broadened its scope in the real world. At the forefront, there is the use under tougher conditions for image recognition. One of these is the application to low-resolution images. In this paper, we discuss how to maximize the estimation accuracy for several typical image recognition problems using images of a lower resolution than those primarily considered in previous studies, as illustrated in the left panel of Fig. 1.
Figure 1: Left: Estimating pose from low-resolution images. Left-Upper: human pose estimation from COCO2017. Left-Lower: head pose estimation from AFLW2000. Right: resolution-accuracy trade-off for human pose estimation evaluated on COCO 2017 val dataset.
@@ I Introduction
Why do we aim to do this? First, multi-scale feature extraction combined with their interactions is an essential component for several tasks. The estimation of human poses mentioned earlier is a prime example of this necessity, as evidenced by previous studies [1, 2, 3]. While the level of necessity may differ, this strategy is fundamentally beneficial across a broad spectrum of image recognition tasks, including, for example, fine-grained object recognition [4] and biomedical image segmentation [5].
@@
Second, low resolution complicates the management of multi-scale features. The conventional approach generates multi-scale features through the downsampling of feature maps [6, 1, 2]. This method is effective when starting with high-resolution images. However, with low-resolution inputs, repeated downsampling quickly diminishes the resolution of feature maps to the point where additional downsampling results in a significant loss of information.
@@
To address this challenge, we investigate methods to manage multi-scale feature maps without downsampling, enabling interactions between features of varying scales. Specifically, we first divide the multi-heads in Multi-Head Self-Attention (MHSA) into several groups, each of which processes features of a different scale. Then, instead of downsampling, we adopt the window-based self-attention mechanism from the Swin Transformer [7] to extract a different scale feature in each group. To facilitate effective interactions among features of different scales, we transfer a feature map from one group to another in an order from lower to higher scales. The transferred feature is utilized in the attention calculations of the subsequent group.
@@
The approach to handling multi-scale features through grouping heads in the first part and the cascading calculation structure in the latter part are inspired by recent CNN-ViT hybrid models, SG-Former [8] and EfficientViT [9], respectively. It is important to note that these existing models, while aiming to address the high computational demands of the ViT architecture and improve the trade-off between performance and computation, were not designed with low-resolution images or the extraction of multi-scale features plus their interactions in mind.
@@
We present experimental results across three distinct tasks: human pose estimation, head pose estimation, and image classification, all utilizing low-resolution image inputs. The results demonstrate that our proposed method achieves superior performance to existing state-of-the-art methods tailored for these individual tasks; see the right panel of Fig. 1. It is also noted that our method needs fewer parameters than the existing methods.
II Related work
II-A Extraction and Use of Multi-Scale Features
@@
Integrating features from multiple scales improves neural networks’ ability to deeply understand complex visual scenes. HRNet [6, 1] and HRFormer [2] fuse the multi-scale representations from multiple branches to effectively generate reliable high-resolution representations with strong position sensitivity. CrossViT [10] learns multi-scale feature representations for image recognition by a dual-branch transformer and merges them multiple times to complement each other. FCT [11] enhances cross-branch interactions in its two-branch transformer model to facilitate multi-scale representation learning. TopFormer [12] produces scale-aware global features from tokens at various scales. P2T [13] and Shunted Transformer [14] improve upon PVT [15] by applying various downsampling rates for the keys and values across heads, rather than using identical spatial reduction for both. This allows each head to capture both coarse- and fine-grained details. SG-Former [8] aggregates tokens at multiple scales and performs attention mechanisms to extract global and fine-grained information for visual tasks. SMT [16] captures diverse spatial features through multiple convolutions with different kernel sizes and conducts scale-aware aggregation to enhance information interaction. SBCFormer [17] achieves efficient global and local feature fusion by introducing a bi-branch CNN-ViT hybrid module. In existing methods, multi-scale representations are typically generated by downsampling the feature map, which can be problematic for low-resolution inputs, as it exacerbates the loss of detail. Our study mitigates this issue by enabling multi-scale feature extraction and interaction without the need for downsampling.
II-B Dealing with Low-Resolution Images
@@
Low-resolution images often lack clarity in fine textures, edges, and small details, presenting significant challenges in image analysis. Some studies tackle this issue by incorporating augmented inputs or utilizing additional knowledge from fully trained teacher models. Haris [18] incorporates high-level vision objectives to improve the model’s capability in detecting low-resolution objects. DeriveNet [19] introduces a multi-resolution pyramid-based data augmentation technique to learn effective features from low-resolution images. NLSA [20] proposes a non-local sparse attention to explore long-range feature correlations in low-resolution images. CAL [21] employs Gaussian offset weighting to account for displacements between coarse predictions and ground-truth positions in low-resolution images. FMD [22] proposes a feature distillation framework to transfer high-definition knowledge from a teacher model for improved low-resolution object recognition. On the other hand, in order to reconstruct fine details from low-resolution images, recent Transformer-based approaches are gaining interest. ESRT [23] combines a CNN backbone with a Transformer backbone to recover a super-resolution image from its degraded low-resolution counterpart. ESSAformer [24] proposes an attention-embedded Transformer to enlarge the receptive field, allowing for better extraction of information from low-resolution hyperspectral images. Steformer [25] designs a bi-branch Transformer-based structure to extract and integrate cross-view information from low-resolution stereo images. However, existing methods do not explore feature fusion in multi-scale space to improve low-resolution image recognition. In this paper, we propose a cascaded multi-scale attention mechanism that constructs enriched features in multi-scale spaces for low-resolution images.
II-C CNN-ViT Hybrids
@@
Hybrid models that combine convolutions and Vision Transformer (ViT) [26] are gaining interest for their ability to capture fine-grained local details while preserving global dependency. To enhance fine-grained information, some methods apply the attention mechanisms to sub-regions or tokens within feature maps. Swin [7] employs shifted windows in the self-attention mechanism to capture both local and global information. CSwin [27] introduces a cross-shaped window to enhance the self-attention mechanism and further improve model capacity. NAT [28] proposes a scalable sliding window attention mechanism by localizing self-attention to the nearest neighboring pixels. QFormer [29] extends the window-based attention mechanism within adaptive quadrangles for learning better feature representation. With the demand for facilitating the aggregation of spatial information, many studies integrate convolutions within the attention mechanism. MobileViT [30] and EfficientFormer [31] combine convolutions with Transformers to achieve efficient local and global feature extraction. IRMB [32] and SBCFormer [17] introduce depth-wise separable convolutions into the self-attention mechanism to improve representational power. FastViT [33] and RepViT [34] employ structural parameterization techniques to optimize computational efficiency and model performance. Recent CNN-ViT hybrid methods demonstrate promising performance in high-resolution image recognition, but their accuracy is limited by the inherent challenges of low-resolution images. In this paper, we propose an effective hybrid network architecture tailored for low-resolution images.
III Attention Mechanism for Low-Resolution Inputs
III-A Motivation
@@
III-B Basic Ideas: Cascaded Multi-Scale Attention (CMSA)
There are two primary ideas. The first involves directing each head of the multi-head self-attention mechanism to handle features from different scales. While conventional multi-head self-attention mechanisms already possess sufficient flexibility, they may not inherently fulfill this role. Contrarily, previous research has shown that attention maps tend to exhibit similarity between heads, leading to redundancy [35, 36].
To address this, we propose organizing the multi-heads into groups and implementing SwinT’s local window attention [7] within each group; see Fig. 2c. We vary the window size for each group. This strategy aims to force different attention heads to effectively manage features of different scales. Our approach draws inspiration from the hybrid-scale attention (HSA) in SG-Former [8]. However, HSA does not use SwinT’s local attention and instead adjusts key and value through token merging [^2] While we initially experimented with the hybrid-scale attention mechanism, we found that our method based on local attention works much better..
Figure 2: Design of the proposed model. (a) The overall architecture. (b) Design of the basic blocks. (c) Proposed cascaded multi-scale attention (CMSA). See texts for details.
Our second idea involves implementing a new mechanism to enhance interaction among features of different scales across various groups. This entails organizing the groups in a specific order and transferring features obtained from local attention within one group to the next. These features are subsequently integrated into the attention computation within each group, resulting in an architecture that operates in a cascade-style fashion, as depicted in Fig. 2c.
This approach draws inspiration from EfficientViT [9]. It divides multi-head attention into groups and performs cascaded information integration, and our method follows the same principle. However, in EfficientViT, processing within groups primarily employs standard self-attention, and the scale of feature maps is the same across all groups. Its aim is to improve the trade-off between computational cost and expressive power. In contrast, our approach emphasizes fostering interactions between features of different scales. We also design a new method for integrating information between groups to achieve this goal.
TABLE I: CMSA variants with different model sizes.
@@ III-B Basic Ideas: Cascaded Multi-Scale Attention (CMSA)
[…]
@@
[…]
@@
headhead=1
[…]
@@
headhead=3
[…]
@@
headhead=4
III-C Details of CMSA
We will now describe the cascaded multi-scale attention mechanism; see Fig. 2c. Let $\mathbf{X}$ be the input to our cascaded multi-scale attention mechanism; $\mathbf{X}$ is a feature map with the size $H\times W\times C$}. We first apply point-wise convolutions (PW-Conv) to $\mathbf{X}$ to obtain $\mathbf{Q}\in\mathbb{R}^{H\times W\times D}$, $\mathbf{K}\in\mathbb{R}^{H\times W\times D}$, and $\mathbf{V}\in\mathbb{R}^{H\times W\times D}$.
@@ III-C Details of CMSA
To deal with multi-scale features, we employ a group structure, as mentioned earlier. Specifically, we divide the heads into $n$ groups, enabling each group to extract and foster features with different scales. Thus, $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$ are segmented in their channel dimensions to yield query, key, and value for each group as $\mathbf{K}_{1}$$\mathbf{K}_{n}$, $\mathbf{Q}_{1}$$\mathbf{Q}_{n}$, and $\mathbf{V}_{1}$$\mathbf{V}_{n}$; $\mathbf{Q}_{k}$, $\mathbf{K}_{k}$, $\mathbf{V}_{k}\in\mathbb{R}^{H\times W\times d_{k}}$.
@@
So far, we have considered the standard setting of attention where the query, key, and value all share the same dimensions. Prior studies [37, 38, 39] demonstrate that making the key and value spatially smaller (or equivalently, using fewer tokens) than the query can enhance computational efficiency with the minimum sacrifice of inference accuracy. Specifically, we adjust the convolution in the spatial fusion (SF) module to halve the spatial dimensions of the key and value compared to the query, while doubling the channel size of the value relative to both the query and key. In our experiments, we employ this method to achieve a better trade-off of accuracy and computational cost.
III-D Design of a Whole Network
III-D1 Overall Design
We adopt a hierarchical pyramidal architecture, in line with recent CNN-ViT hybrids; see Fig. 2a. The architecture starts with a ‘Stem’ block, followed by a sequence of stages. Each stage comprises multiple blocks that include the cascaded multi-scale attention mechanism explained above. A patch embedding layer is placed between stages to reduce the spatial resolution of feature maps.
@@ III-D1 Overall Design
III-D2 Design of a Block
As shown in Fig. 2b, a single block consists of three components: position embedding, the cascaded multi-scale attention, and a feed-forward network.